OpenStudy (anonymous):

A. 8 and 9 B. 9 and 10 C. 10 and 11 D. 84 and 85

OpenStudy (anonymous):

i will fan and give a medal

sam (.sam.):

First you have to find what number when squared is the narrowest range of \(\sqrt{85}\)

sam (.sam.):

For this, try \[\sqrt{84}\] That doesn't give an integer try \[\sqrt{83}\] That doesn't give as well, how about \[\sqrt{81}\] This is equal to 9, so 9 will be lower limit of \(\sqrt{84}\)

sam (.sam.):

The upper boundary you'll need to test is \(\sqrt{86},\sqrt{87},\sqrt{88},\sqrt{89}...\) Try until you got an integer

sam (.sam.):

I think the closest you can get is \(\sqrt{100}\) which is 10, so the answer is B

OpenStudy (anonymous):

What is the value of x?